11. The demand for a monopolists product is given by Q-400 4P while the monopolists marginal cost is given by MC- Q The profit-maximizing quantity of output for this monopolist is A) 1o0 B) 44-44 D) 20 There is a payoff matrix of two firms; their different profits are listed when they choose collusion or competition (answer 14-15). firm B competition 4.14 5. 28 collusion 27,5 19, 19 firm A competition collusion 12. In the game above, who has dominant strategy? A) Player A has a dominant strategy C) Both players have dominant strategies. D) Neither player has a dominant strategy B) Player B has a dominant strategy 13. Which one is the Nash equilibrium? A) firm A competition, firm B competition. B) firm A competition, firm B collusion. C) firm A collusion, firm B competition. D) firm A collusion, firm B collusion. In the following game, all payoffs are listed with the row players payoffs first and the column players payoffs second (a9-20). Player B B1 B2 Player A Al 5,6 7,2 A2 4,5 9, 1 14. In game above, a) Player A choosing Ai and Player B choosing Bi is a Nash equilibrium. b) Player A choosing A2 and Player B choosing Bz is a Nash equilibrium. c) there is no Nash equilibrium. d) there are multiple Nash equilibria in pure strategies. 15. In game above, a) Player A has a dominant strategy b) Player B has a dominant strategy c) Both players have dominant strategies. d) Neither player has a dominant strategy Part 2: Short Questions (You need to show steps to get full credits) 1. Meimei Hans utility function is U-ioXZ, The price of X is sio, the price of Z is
Answer #1

11. The demand function is Q = 400 - 4P, Marginal Cost: MC = 2Q

The monopolist's profit is maximised at the point where Marginal Revenue (MR) = Marginal Cost (MC)

Now, deriving P from the demand function we get,

Q = 400 - 4P

4P = 400 - Q

P = (400 - Q)/4

Total Revenue (TR) = Price * Quantity

TR = [(400 - Q)/4]Q

= (400Q - Q2)/4

Marginal Revenue (MR) = \partial TR/\partial Q = (400 - 2Q)/4

Since, the profit maximizing quantity of output for this monopolist is Marginal Revenue = Marginal Cost

we will equate MR with MC to get the profit maximizing quantity of output.

(400 - 2Q)/4 = 2Q

400 - 2Q = 8Q

10Q = 400

Q = 40

which is Option C) 40